# IEEEexample:BSTcontrol

### v1

@@
Abstract
The advent of internet medicine provides patients with unprecedented convenience in searching and communicating with doctors relevant to their diseases and desired treatments online. However, the current doctor recommendation systems fail to fully ensure the professionalism and interpretability of the recommended results. In this work, we formulate doctor recommendation as a ranking task and develop a large language model (LLM)-based pointwise ranking framework. Our framework ranks doctors according to their relevance regarding specific diseases-treatment pairs in a zero-shot setting. The advantage of our framework lies in its ability to generate precise and explainable doctor ranking results. Additionally, we construct DrRank, a new expertise-driven doctor ranking dataset comprising over 38 disease-treatment pairs. Experiment results on the DrRank dataset demonstrate that our framework significantly outperforms the strongest cross-encoder baseline, achieving a notable gain of +5.45 in the NDCG@10 score while maintaining affordable latency consumption. Furthermore, we comprehensively present the fairness analysis results of our framework from three perspectives of different diseases, patient gender, and geographical regions. Meanwhile, the interpretability of our framework is rigorously verified by three human experts, providing further evidence of the reliability of our proposed framework for doctor recommendation.
{IEEEkeywords}
Doctor recommendation, explainable recommendation, large language model, text ranking.
@@ 1 Introduction
[…]
The rapid development of internet medicine provide patients with convenient access to trans-regional disease diagnosis and treatment services [1]. Through professional online healthcare platforms, patients can search for and select experienced doctors based on their potential diseases and required treatments, including medication, surgery, and radiotherapy. Although some authoritative leaderboards have been released, such as the China Hospital and Departments Reputation Ranking [^1] https://rank.cn-healthcare.com/fudan/national-general, this leaderboard can only guide patients in choosing from a limited number of hospitals and departments, let alone specific doctors. For patients lacking medical knowledge, accurately selecting a doctor who matches their condition remains challenging. To provide high-quality online medical services, a crucial task is to recommend an appropriate doctor from a large pool of candidates for a specific patient [2, 3].
Recent studies have shown significant interest in developing doctor recommendation systems. Early methods rely on basic statistical indicators such as the number of patient reviews and treatment cases [4, 5]. These methods fail to fully capture the nuanced professional expertise of doctors, consequently compromising the accuracy of rankings. Current methods seek to improve doctor recommendation through machine learning techniques, but some significant limitations remain. Machine learning methods usually require sufficient labeled data for model training, which poses a significant challenge in data-sparse medical scenarios due to the high cost and effort of data annotation. Some studies define the ground truth label in doctor-patient interactions based on whether the consultation order is successfully completed by both parties [6], or whether the doctor later responded to the patient’s inquiry [7]. In such cases, the optimization objectives of recommendation systems tend to be skewed towards commercial considerations. This commercial bias, to some extent, undermines the focus on doctors’ expertise and may not adequately address patients’ needs to find experts in specific medical fields. Additionally, explanations are crucial in ensuring the interpretability of recommendation systems [8], particularly in the medical context of doctor recommendation. However, the interpretability of doctor recommendation is still largely unexplored at present.
Figure 1: An overview of our proposed zero-shot LLM-based doctor ranking framework. We leverage a pointwise ranking method to efficiently and effectively evaluate the professional relevance of candidate doctors in addressing specific patient medical needs. By utilizing fine-grained relevance labels and incorporating query-specific ranking criteria, our framework ensures consistent and accurate doctor evaluations. Furthermore, it provides clear and structured doctor evaluation rationales as recommendation explanations, enhancing the interpretability and transparency of the doctor recommendation process. It should be noted that the ranking criteria and recommendation explanation shown in the figure are simplified versions.
In this work, we focus on constructing an accurate, fair, and explainable doctor recommendation system based on the professional relevance of doctors in addressing patient medical needs. We formulate doctor recommendation as a ranking task and introduce DrRank, a novel expertise-driven doctor ranking dataset that spans various disease-treatment pairs. To create comprehensive doctor profiles, we integrate multidimensional professional information, such as their introductions, areas of expertise, community involvement, awards received, and research achievements. Furthermore, we involve medical experts to annotate the professional relevance of doctors in relation to specific disease-treatment pairs. In the field of information retrieval, cross-encoder models rely on annotated relevance data for supervised training [9]. These models have exhibited impressive performance in various out-of-domain ranking tasks [10]. However, their efficacy in the specific context of doctor ranking has yet to be explored, primarily because such a task is typically absent from previous training datasets.
@@
Large language models (LLMs), backed by extensive pre-training data, demonstrate the ability to perform text ranking tasks based on user prompts [11, 12, 13, 14]. The remarkable generalization capabilities and inherent world knowledge of LLMs suggest significant potential for the doctor ranking task. Inspired by this, we propose an LLM-based doctor ranking framework as illustrated in Fig. 1, which ranks doctors according to their relevance regarding specific disease-treatment pairs. It operates in a zero-shot setting, thus eliminating the need for vast amounts of labeled data for training Specifically, we employ a pointwise ranking method to ensure the efficiency of the doctor ranking process. We utilize fine-grained relevance labels to assist the LLM in more accurately distinguishing the nuanced differences in doctors’ relevance across various disease-treatment pairs. To enhance the consistency and accuracy of the LLM’s ranking results, we prompt the LLM to generate specific ranking criteria tailored to a specific disease and treatment option. The generated ranking criteria serve as evaluation guidelines, providing clear and transparent directives for the LLM during the ranking decision-making process. To improve the interpretability of the doctor recommendation process, once the LLM has generated doctor relevance predictions, we continuously prompt it to further generate evaluation rationales as the explanations for doctor recommendation. We observe that the stochastic behavior of the LLM can potentially influence both the relevance generation and rational generation stages. To address this, we propose a forced elicitation mechanism that ensures the LLM consistently generates the desired results.
Experimental results demonstrate that LLMs, with a scale exceeding 7 billion parameters, exhibit superior performance compared to cross-encoder models for doctor ranking task, particularly when ranking criteria are introduced. Additionally, we present the time costs associated with utilizing the LLM for pointwise prediction and the generation of evaluation rationales, illustrating the feasibility of our framework in real-world applications. Furthermore, we analyze the distribution characteristics of the LLM prediction results and ground truth relevance labels, revealing how our framework works in practice. Meanwhile, we evaluate the fairness and interpretability of our framework. Experimental results show that our framework exhibits better fairness performance compared to cross-encoder models across a range of disease-treatment pairs. However, we also acknowledge concerns regarding potential gender biases. Through manual evaluation of the ranking criteria and evaluation rationales within our framework, we highlight the significant potential and practical utility of our framework in serving patients.
@@
  1. 1.
    We construct a new expertise-driven dataset for objectively evaluating models in doctor ranking.
  2. 2.
    We develop a zero-shot LLM-based doctor ranking framework for doctor recommendation, addressing the issue of limited labeled data for model training.
  3. 3.
    Through extensive experiments, we demonstrated the accuracy, fairness, and interpretability of our framework, and we have released our code to ensure reproducibility.
@@ 2.1 Doctor Recommendation
Doctor recommendation aims to help patients find the most suitable doctors. While many works in recommendation focus on modeling target users from historical behavior [15], doctor recommendation can only rely on limited information in patients’ chief complaints for modeling due to privacy concerns. Modeling doctors effectively is another pivotal area of research. Some studies focus on constructing doctor profiles based on basic features available on online healthcare platforms [16, 17], such as years of experience, pricing, and online reviews. Furthermore, dialogue data between doctors and patients has been utilized to enhance doctor modeling and to investigate the combined effects of doctor profiles and previous dialogues with other patients [6, 7]. Additional research efforts are directed towards developing personalized recommendation to help patients find doctors who align with their personal preferences [18, 19]. Che et al. [20] delve into the challenge of integrating data of varying structures across multiple medical platforms, exploring innovative ways to leverage this diverse information effectively. In this work, we focus on doctor recommendation based on comprehensive doctor profiles and leverage the zero-shot ranking capabilities of LLMs to escape the predicament of training a doctor ranking model with scarce labeled data.
@@ 2.2 Large Language Models in Zero-shot Text Ranking
Shifting from tuning-based learning to rank driven by cross-encoder models [9, 10] such as MonoT5 [21] and RankT5 [22], there is an emerging line of research exploring how to utilize general-purpose LLMs for zero-shot text ranking in out-of-domain scenarios. The current LLM-based ranking methods can be mainly divided into the following three categories: (1) Pointwise strategy utilizes LLMs to measure the relevance between a query and a single passage by query generation [23] or relevance generation [24, 11]. (2) Pairwise strategy prompts LLMs to determine which of two passages is more relevant to a query and aggregates a final relevance score for each passage from all pair comparisons, where efficient sorting algorithms, such as heapsort, can be used to speed up the ranking process particularly in top-k ranking [13]. Setwise strategy enhances the efficiency of pairwise comparisons by selecting the most relevant passage from multiple candidates [14]. (3) Listwise strategy instructs LLMs to generate a permutation of passages, arranged in descending order of relevance to a query [12]. Due to the limited input length of LLMs, a sliding window approach is used to rerank a subset of candidate passages each time, moving the window from back to front to cover all candidates. In this work, we formulate doctor recommendation as a ranking task and develop a zero-shot LLM-based pointwise ranking framework which focuses on the accuracy, consistency and interpretability of doctor ranking.
@@ 3 DrRank Dataset
Figure 2: An example of doctor profiles used in the DrRank dataset, highlighting the nine key fields. This detailed profile format ensures comprehensive representation of a doctor’s professional background and expertise.
@@ 3.1 Data Collection and Preprocessing
The doctor profiles used in this study is primarily sourced from Haodf Online [^2] https://www.haodf.com, one of the biggest online healthcare platforms in China. Each doctor profile comprises 9 fields: “Title”, “Specialty”, “Affiliation”, “Department”, “Introduction”, “Expertise”, “Social Service”, “Awards”, and “Research”, as depicted in Fig. 2. Over 600,000 doctor profiles were collected in total, forming a comprehensive resource for analysis. To proceed, we first selected a range of common cancers and critical illnesses, along with their typical treatment options, resulting a set of 38 disease-treatment pairs. For each pair, such as lung cancer and surgical treatment, we employed the open-sourced Alibaba-NLP/gte-multilingual-reranker-base [^3] https://huggingface.co/Alibaba-NLP/gte-multilingual-reranker-base model [25] to assess the relevance of the introductions from the doctor profiles in our extensive pool to the disease. This process yielded an initial list of 100 candidate doctors for each disease-treatment pair.
Figure 3: Cumulative distribution functions (CDF) of doctor profile lengths from two datasets: DrRank and Collected Doctor Pool. The x-axis represents the length of tokenized doctor profiles, and the y-axis represents the cumulative percentage of sequences. The vertical dashed lines at 1024, 2048, and 4096 tokens indicate key truncation thresholds. Each of these thresholds is annotated with the percentage of the DrRank dataset that is covered up to that token count.
@@ 3.2 Data Annotation
For each disease-treatment pair, two experts in this field were interviewed for approximately one hour. They evaluated the relevance of the doctors on our initial list and also supplemented it with additional professional doctors who were missing from the list but should be included based on the actual situation. Each doctor on the list was assigned a professional relevance score ranging from 0 to 5, with 5 indicating the highest relevance and 0 representing irrelevance. Doctors who received consistent relevance scores from both experts were retained directly. In cases of disagreement between the two experts regarding the assigned relevance score, a third independent expert was consulted to provide the final evaluation.
@@ 3.3 Hard Negative Mining
We considered doctors with a relevance score greater than 1 as positive samples and those with a relevance score of 0 as negative samples. To balance the number of positive and negative samples for each disease-treatment pair and enhance the complexity of distinguishing between positive and negative samples, we conducted hard negative mining. Through statistical analysis of the doctor profiles of annotated positive samples, we found that they tend to be relatively long. Therefore, we filtered the doctor profiles to include only those with introductions exceeding a minimum length of 1,024  [^4] Unless otherwise specified, all lengths mentioned in this work are defined based on the tokenizer inherent to the Qwen2.5 series.. This filtering process resulted in a subset of 5,000 unique doctors with sufficiently rich profile information. For each disease-treatment pair, we utilized the gte-multilingual-reranker-base [25] model once again to rank the filtered set of doctors based on the relevance score between the disease and the doctor’s introduction. To minimize false negatives, we excluded the top 1% of doctors from the ranked list. Then, from the remaining list, we iteratively added negative samples not initially included, in descending order of relevance, until we achieved a 1:1 ratio of positive to negative samples for each disease-treatment pair. To further introduce more challenging negative samples, we randomly replaced 30% of the previously selected negative samples with positive samples from other pairs. Specifically, we regarded the positive samples with relevance scores of 4 and 5 from other disease-treatment pairs as negative samples for the current pair. Finally, we conducted a manual review to ensure that the all selected negative samples for each disease-treatment pair were indeed unrelated to it.
Figure 4: Heatmap of the annotated professional relevance distribution across positive samples for different disease-treatment pairs in the DrRank dataset. The x-axis represents the professional relevance (ranging from 1 to 5), while the y-axis lists various disease-treatment pairs.

### v2

# 5- BIG DATA MINING AND ANALYTICS

@@
[…]
-5Big Data Mining and Analytics, January 2018, 1(1): $000\headoddnameZeng$ et al.:A Zero-shot Explainable Doctor Ranking Framework with Large Language Models
[…]
5- BIG DATA MINING AND ANALYTICS
DOI:  1 0 . 2 6 5 9 9 / B D M A . 2 0 2 5 . 9 0 2 0 0 9 8
[…]
5- Volume  1,  Number  1,  January  2018
[…]
5- V o l u m e 1, N u m b e r 1, S e p t e m b e lr 2 0 1 8
$\bullet$ Ziyang Zeng and Yuqing Yang are with School of Information and Communication Engineering, Beijing University of Posts and Telecommunications, Beijing 100876, China. Email: ziyang1060@bupt.edu.cn; yangyuqing@bupt.edu.cn
$\bullet$ Dongyuan Li is with School of Artificial Intelligence, Beijing University of Posts and Telecommunications, Beijing 100876, China. Email: li_dongyuan@bupt.edu.cn
[…]
Manuscript received: 2025-Apr-19; accepted: 2025-Sep-08
[…]
5
@@ 1 Introduction
The rapid development of internet medicine has greatly enhanced patients’ access to trans-regional diagnostic and treatment services [1]. Through professional online healthcare platforms, patients can search for and choose experienced doctors based on their suspected conditions and treatment needs, such as medication, surgery, or radiotherapy. However, for patients with limited medical knowledge, selecting an appropriate doctor from a vast pool of professionals remains a significant challenge.
To improve the quality of online healthcare service, a key task is doctor ranking—that is, ranking doctors from a large candidate pool based on their relevance to a patient’s medical needs [2, 3]. This task has attracted growing attention in recent studies. Early approaches typically rely on simple statistical indicators, such as the number of consultations or patient reviews [4, 5]. However, these metrics do not adequately reflect a doctor’s professional expertise, often resulting in suboptimal rankings. More recent methods adopt machine learning techniques to improve ranking accuracy [6, 7]. Nevertheless, these models generally depend on labeled data for supervised learning, which is challenging to obtain in the medical domain due to the high cost of expert annotation. Some studies attempt to construct supervised signals using proxy objectives—such as whether a consultation was completed [8] or whether the doctor responded to an inquiry [9]. However, such heuristic signals often introduce commercial bias, shifting focus away from professional competence and compromising the reliability of ranking results for patients seeking domain-specific expertise. Despite these advancements, existing methods largely overlook the explainability of doctor ranking, which is critical for building trustworthy recommendation systems [10]. Meanwhile, some commercial medical platforms—for example Haodf Online***https://www.haodf.com—have released department-specific doctor rankings, but without disclosing the internal mechanisms behind the rankings. In high-stakes domains such as healthcare, providing transparent explanations is indispensable to foster patient trust and secure acceptance of doctor rankings.
Fig.  1: An overview of our proposed zero-shot, large language model-based, explainable doctor ranking framework.
@@
Large language models (LLMs), backed by extensive pre-training on large-scale textual data, have demonstrated impressive capabilities in performing text ranking tasks based on user-defined prompts [11, 12, 13, 14]. Their remarkable generalization abilities and embedded medical knowledge make them highly suitable for the doctor ranking task, especially in data-scarce environments. Inspired by this, we propose a novel large language model (LLM)-based doctor ranking framework, as illustrated in Fig. 1, which ranks doctors according to their relevance to specific disease-treatment pairs. The framework operates in a zero-shot setting, thereby eliminating the need for large-scale labeled training data. To ensure efficiency, we adopt a pointwise ranking strategy, where the LLM predicts fine-grained relevance labels for individual doctor profiles. These precise relevance labels help the model distinguish nuanced differences in professional expertise across a variety of disease-treatment scenarios. To further enhance the consistency and interpretability of doctor ranking, we prompt the LLM to generate disease-specific ranking criteria, which serve as structured guidelines for relevance evaluation during the ranking decision-making process. After generating relevance labels, the LLM is further prompted to produce step-by-step rationales, which serve as explanations for its ranking decisions. We observe that the stochastic behavior of LLMs may affect both the relevance estimation and rationale generation. To address this, we design a forced elicitation mechanism that encourages the LLM to consistently produce both relevance scores and their corresponding justifications in a controlled manner.
To enable rigorous and objective evaluation of doctor ranking, we introduce DrRank, the first expertise-driven doctor ranking benchmark dataset covering 38 disease-treatment pairs and 4,325 doctor profiles. Each doctor profile is constructed using rich multidimensional information, including personal introductions, areas of expertise, community engagement, honors and awards, and academic achievements. Moreover, we collaborate with certified medical experts to annotate doctor relevance in relation to specific disease-treatment pairs, providing high-quality supervision for evaluation.
Experimental results demonstrate that LLMs with more than 7 billion parameters significantly outperform baseline models, especially when guided by the generated ranking criteria. We further analyze the distribution characteristics of relevance scores with respect to the ground-truth label, revealing how our framework operates in practice. In addition, we evaluate the fairness of our framework, with results showing a significant advantage in fairness across disease types and patient gender over baseline methods. Furthermore, through manual evaluation of the generated criteria and rationales, we highlight the framework’s practical utility and trustworthiness, reinforcing its potential for real-world doctor recommendation. This strong performance is not limited to the medical domain; our framework’s robustness and generalization are also validated on two datasets from BEIR benchmark, where it again achieves impressive zero-shot performance, confirming its broader applicability as a general-purpose ranking framework.
@@
  1. (1)
    We propose a zero-shot, LLM-based explainable doctor ranking framework that effectively addresses the challenge of limited labeled data in medical domains.
  2. (2)
    We construct and release DrRank, the first expertise-driven benchmark dataset for objective evaluation of doctor ranking systems.
  3. (3)
    Through extensive experiments, we demonstrated the accuracy, fairness, interpretability and generalizability of our framework, and we have released our code to ensure reproducibility.
@@ 2.1 Doctor Recommendation
Doctor recommendation aims to assist patients in identifying the most suitable medical professionals. Unlike conventional recommendation systems that model user preferences based on historical behavior [15], doctor recommendation and ranking often rely on limited patient information—typically derived from brief chief complaints—due to privacy concerns. Doctor modeling is another pivotal research focus in this domain. Early approaches construct doctor profiles based on readily available attributes from online healthcare platforms [6, 7], such as years of experience, consultation fees, and user reviews. More recent studies leverage dialogue data between doctors and patients to enhance doctor representations and investigate how past interactions influence model predictions [8, 9]. Other works explore personalized recommendation strategies to better align patients with doctors based on individual preferences and healthcare needs [16, 17]. In addition, Che et al. [18] address the integration of heterogeneous data across multiple healthcare platforms, proposing methods to unify and utilize diverse sources of information.
Doctor ranking is typically considered a component of broader doctor recommendation systems. In this work, we focus exclusively on this aspect, aiming to generate accurate and explainable rankings that can serve as a foundation for building trustworthy recommendation systems in real-world healthcare platforms. Yet, the absence of a standardized benchmark in this field makes it difficult to conduct fair comparisons across different approaches. To address this gap, we present a professionally annotated doctor ranking dataset for the research community and benchmark general-purpose ranking models [19]—including cross-encoder and bi-encoder architectures—as zero-shot baselines for future research.
@@ 2.2 Large Language Models in Zero-shot Text Ranking
Shifting from tuning-based learning to rank driven by pretrained models [19] such as MonoT5 [20] and RankT5 [21], there is an emerging line of research exploring how to utilize general-purpose LLMs for zero-shot text ranking in out-of-domain scenarios. The current LLM-based ranking methods can be mainly divided into the following three categories: (1) Pointwise strategy utilizes LLMs to measure the relevance between a query and a single passage by query generation [22] or relevance generation [23, 11]. (2) Pairwise strategy prompts LLMs to determine which of two passages is more relevant to a query and aggregates a final relevance score for each passage from all pair comparisons, where efficient sorting algorithms, such as heapsort, can be used to speed up the ranking process particularly in top-k ranking [13]. Setwise strategy enhances the efficiency of pairwise comparisons by selecting the most relevant passage from multiple candidates [14]. (3) Listwise strategy instructs LLMs to generate a permutation of passages, arranged in descending order of relevance to a query [12]. Due to the limited input length of LLMs, a sliding window approach is used to rerank a subset of candidate passages each time, moving the window from back to front to cover all candidates. In this work, we leverage the zero-shot pointwise ranking capabilities of LLMs to address the challenge of training doctor ranking models with limited labeled data.
2.3 Explainable Ranking
In recent years, the field of explainable ranking has garnered significant attention, aiming to enhance the transparency and trustworthiness of ranking models. Existing approaches can be broadly categorized into two lines of work: intrinsically interpretable ranking models and natural language-based explanations. The first line focuses on designing models that are inherently interpretable, such as additive models [24], which provide clear insights into how individual human-engineered ranking features [25] contribute to ranking decisions. However, as ranking tasks grow increasingly complex, the performance of such interpretable models is often limited by their simplistic architectures. A more promising direction involves leveraging natural language generation to produce human-understandable explanations for ranking outcomes. For example, ExaRanker [26] incorporates synthetic explanations generated by LLMs to enhance both training efficiency and the interpretability of sequence-to-sequence ranking models. LiEGe [27] jointly explains all documents in a ranked list by learning semantic representations at two levels of granularity—documents and their constituent tokens. A work most related to ours is MCRanker [28], which enhances LLM-based pointwise ranking through multi-perspective criteria. This approach can be viewed as a more complex variant of ours, involving multiple rounds of LLM inference and result ensembling from different perspectives. While it potentially offers improved performance, it significantly increases computational overhead and system complexity.
However, prior works typically focus on either the ranking process or the ranking output, and rarely integrate both into a unified explainable framework. In contrast, our approach jointly models ranking and explanation generation, using LLMs to produce both explicit criteria and step-by-step rationales. This unified framework simultaneously enhances the interpretability of both the decision-making process and the final outcomes, making the overall pipeline more transparent and trustworthy.
@@ 3 DrRank Dataset
Fig.  2: An example of doctor profiles used in the DrRank dataset.
@@ 3.1 Data Collection and Preprocessing
The doctor profiles used in this study is primarily sourced from Haodf Online, one of the biggest online healthcare platforms in China. Each doctor profile comprises 9 fields: “Title”, “Specialty”, “Affiliation”, “Department”, “Introduction”, “Expertise”, “Social Service”, “Awards”, and “Research”, as depicted in Fig. 2. Over 600,000 doctor profiles were collected in total, forming a comprehensive resource for analysis. To proceed, we first selected a range of common cancers and critical illnesses, along with their typical treatment options, resulting a set of 38 disease-treatment pairs. For each pair, such as lung cancer and surgical treatment, we employed the open-sourced gte-multilingual-reranker-basehttps://huggingface.co/Alibaba-NLP/gte-multilingual-reranker-base model [29] to assess the relevance of the introductions from the doctor profiles in our extensive pool to the disease. This process yielded an initial list of 100 candidate doctors for each disease-treatment pair.
@@ 3.2 Data Annotation
The annotation team was composed of clinical experts from Peking Union Medical College Hospital—consistently ranked as China’s top hospital in comprehensive evaluations. For each disease–treatment pair, two specialists in the relevant field were interviewed for approximately one hour. During these interviews, they reviewed our preliminary list of doctors, assessed each individual’s suitability, and recommended any additional professionals who, based on real-world practice, should also be included. Evaluation criteria encompassed disease relevance, hospital tier, physician title, positions held within professional societies, research accomplishments, and honors received. Each doctor was then assigned a relevance score from 0 to 5 (5 = highest relevance; 0 = irrelevant). In practice, these scores correspond to six tiers: leading experts, nationally recognized specialists, provincially acknowledged experts, municipal-level authorities, general specialists, and irrelevant physicians. Doctors whose scores agreed between both reviewers were retained outright; whenever the two experts’ scores diverged, a third independent specialist adjudicated to determine the final score. It is worth noting that the annotated dataset encompasses professionals from a wide range of institutions rather than focusing solely on top-tier hospitals.
Fig.  3: Heatmap of the annotated professional relevance distribution across positive samples for different disease-treatment pairs in the DrRank dataset.
@@ 3.3 Hard Negative Mining
To balance the number of positive and negative samples for each disease-treatment pair and enhance the complexity of dataset, we conducted hard negative mining. We considered doctors with a relevance score greater than 1 as positive samples and those with a relevance score of 0 as negative samples. Statistical analysis of the doctor profiles in the annotated positive samples revealed that they tend to be relatively long. Therefore, we filtered the profiles to include only those with a total length exceeding 1,024 Unless otherwise specified, all lengths mentioned in this work are defined based on the tokenizer inherent to the Qwen2.5 series.. This filtering process resulted in a subset of 5,000 unique doctors with sufficiently rich profile information. For each disease-treatment pair, we utilized the gte-multilingual-reranker-base again to rank the filtered set of doctors based on the relevance score between the disease and the doctor’s introduction. To minimize false negatives, we excluded the top 1% of doctors from the ranked list. Then, from the remaining list, we iteratively added negative samples not initially included, in descending order of reranker scores, until we achieved a 1:1 ratio of positive to negative samples for each disease-treatment pair. To further introduce more challenging negative samples, we randomly replaced 30% of the previously selected negative samples with positive samples from other pairs. Specifically, we regarded the positive samples with relevance scores of 4 and 5 from other disease-treatment pairs as negative samples for the current pair. Finally, we conducted a manual review to ensure that the all selected negative samples for each disease-treatment pair were indeed unrelated to it.